# StoX-Net: Stochastic Processing of Partial Sums for Efficient In-Memory Computing DNN Accelerators

### v1

@@ 3.2. Hardware-Aware Training of Quantized DNN with Stochastic Partial Sum
Where is the layer’s final output obtained from the summation of all array-level output (each sample and each subarray ). For each subarray output vector and its respective MTJ samples, the gradient needs to be calculated as:
Equation 3 considers the , MTJ, weight quantization, array-splitting, and normalization algorithms with respect to the layer’s weights. We approximate the MTJ’s gradient using a straight-through estimator (STE), and clamp values outside its saturation range to avoid exploding gradients. The weight quantization’s gradients are similar to that of (Qin et al., 2020). Considering all of these, the gradient for backward propagation can be reduced to:
as shown in Fig. 7.
4. Experimental Results: Functional accuracy and Hardware efficiency
4.1. Evaluation Methodology
We evaluate both the functional accuracy and hardware efficiency of the proposed StoX-Net designs, including both a quantized first layer (QF) and a high-precision first layer (HPF), while other subsequent layers adopt the MTJ-based stochastic conversion. All QF models take 8 samples per MTJ conversion in the first layer due to its importance shown by the Monte Carlo analysis in Fig. 6. We compare StoX-Net with a standard IMC featuring a high-precision ADC for all layers (HPFA). We also include an IMC design with sparse ADC (SFA) as a strong baseline for hardware comparison, where we reduce the precision of the ADC by 1 from the full precision ADC in HPFA. We investigate the impact of hardware and software configurations, including array sizes, weight bit precision, tanh sensitivity parameter, and the number of MTJ samples across the study.
We denote X-bit weight, Y-bit activation, and Z-bits per slice as . For example, 4-bit weight, 4-bit activation, and 2-bits per slice are . Unless otherwise specified, the baseline StoX network can be assumed, having characteristics of , , , 1 sample per MTJ, and HPF. Networks will only be described by their difference from this baseline. For example, 1-QF uses the stochastic MTJ converter with 1 sample in all layers except the first, which would have an MTJ converter with 8 samples. 4-HPF uses a high-precision ADC in the first layer and 4 samples of the stochastic converter in all other layers. Layers with high- will not see multisampling as such hardware is deterministically binary.
@@ 4.1. Evaluation Methodology
For StoX-Net’s accuracy measurements, we train Resnet-20 models based on the proposed convolutional methodology in Section 3. For CIFAR-10, we train with hyperparameters similar to (Qin et al., 2020), and for MNIST we lower epochs from 400 to 25.
For evaluating hardware efficiency, we simulate a simple IMC architecture similar to that of ISAAC (Shafiee et al., 2016), as shown in Fig. 8. To account for both positive and negative input activations and weights, we follow the representation described in (Saxena et al., 2022). In this representation, the positive and negative components of the activations are streamed separately. Two bits corresponding to the positive and negative components of each weight bit are stored in a group of two memristor cells. As a result, the current accumulated in each crossbar column, which can be either positive or negative, represents the MVM operation. Hence, the energy and area values for the DAC and crossbar (Xbar) cells are doubled during the hardware evaluation. We create our architecture models using Accelergy/Timeloop (Parashar et al., 2019; Wu et al., 2019, 2020). We use ReRAM devices with () of 500kOhm, and an ratio of 10 (Ankit et al., 2019) as the Xbar cells. We use Successive Approximation Register (SAR) ADC with the area and power metrics scaled for 28nm technology based on comprehensive surveys and analysis of ADC architectures (Murmann, [n. d.]) (He et al., 2022; Wang et al., 2016). Table 2 summarizes key parameters for the hardware components. Our baseline HPFA architecture results were verified to be consistent with previous works (He et al., 2022)(Ankit et al., 2019).
4.2. Functional Accuracy
| StoX-Net | Ref.(Saxena et al., 2022) | |||
| Samples: | 1-QF | 4-QF | Mix-QF | HPF |
@@ 4.2. Functional Accuracy
| StoX-Net | Ref. (Saxena et al., 2022) | Ref. (Qin et al., 2020) | ||||
| 1 | 4 | 8 | Mix | High- HPF | ||
| QF | 76.1 | 83.3 | 83.8 | 80.6 | – | – |
| HPF | 84.3 | 86.0 | 86.6 | 85.5 | 84.2 | 85.4 |
As summarized in Tables 3 & 4, the proposed StoX-Net can reach within 1% of the reference HPF model’s accuracy for both MNIST and CIFAR-10 datasets. As shown in Fig. 9A, large arrays suffer small accuracy loss as the stochastic conversion is applied to an increased range of possible crossbar output values, losing information. Bit slicing is shown to moderately mitigate the accuracy degradation of quantization (Fig. 9B). Most notably, stochastic multisampling consistently mitigates accuracy loss.
In Fig. 9E, ”high-, high- QF” implements step-like sense amplifier to quantize all layers including the first conv layer. Comparatively, by enabling a stochastic first layer, ”high-, QF” experiences immediate (over 3.4%) accuracy improvement, with 8 samples for the first StoX conv (see the bright-green and grey bars). Such observation demonstrates that our stochastic MTJ enables an effective quantization of the first conv layer through its increased representational capability.
We also observe that while larger (step-like) might be preferred for 1-time sampling (Fig. 9D), higher accuracy can be obtained through multisampling (Table 4), such as 4 or 8 samples, of a smaller (stochastic). The configuration of layer-wise mixed samplings (”Mix-QF”) only increases the number of MVM conversions by 14.3% compared to the 1-sample model but reaches 85.5% accuracy near the 4-sample network. As discussed in the following section, although sampling multiple times monotonically increases the energy and latency, our 4-QF configuration that uses 4 samples per MTJ conversion still achieves 100x improvement in EDP.
4.3. Hardware Efficiency
@@
It is important to note that our area-efficient and highly parallel PS processing directly impacts the datapath design pattern of crossbar MVM operation. Fig. 10 shows the pipelined operations inside a crossbar using an ADC (top) and a stochastic MTJ converter (bottom). In standard IMC, the length of each pipeline stage is determined by the longest stage, i.e. the ADC readout of all shared columns in a crossbar. Our optimization alleviates the throughput bottleneck by parallelizing all columns with compact MTJ converters, thereby considerably reducing the length of the pipeline stage (from 128 ns to 1.85 ns in our case). However, increasing the number of samples reduces such throughput benefits and creates a trade-off between latency and accuracy.
@@
Our hardware performance evaluation is summarized in Fig. 11. Full precision ADC (HPFA) and Sparse design (SFA) are our baselines for comparison. The first observation is that SOT-MTJ converters significantly enhance parallelism without large area overhead, showing large savings in area and EDP. Our proposed hardware design can offer area, energy, and latency reduction of up to , , and , respectively. Our design achieves up to () EDP improvement compared to an IMC with sparse (full-precision) ADC. We further demonstrate in Fig. 11(b) that our architecture can be scaled to larger models and workloads while maintaining remarkable efficiency improvement. Moreover, the multisampling feature of our proposed stochastic converter offers the flexibility to trade off EDP gain in exchange for improved accuracy.
5. Conclusion
@@
We develop StoX-Net, an efficient IMC design that eliminates the ADC bottleneck by processing array-level partial sum using probabilistic switching of MTJ. The proposed IMC co-design framework reaches close to the state-of-the-art accuracy at CIFAR-10 classifications while achieving over 140x area efficiency and 100-600x improvement in hardware efficiency characterized by EDP. Moreover, leveraging the multisampling capability of stochastic MTJ converters, quantizing the first convolution layer, which is costly in prior quantization-aware training, was demonstrated to have less than accuracy loss. Our work suggests exciting opportunities for utilizing stochastic computation and spintronics for developing next-generation AI hardware accelerators.

### v2

@@ 3.2.2 Backpropagation with Stochastic binarization
Where is the layer’s final output summing all the array-level outputs after the shift-and-add operation (). For each layer output vector that goes through and sampling through the MTJ converter, the gradient is calculated as follows:
Each array ’s gradient backtracks to the shift-and-added (S&A) result of the stochastic MTJ output, which counters (sums) the output over MTJ samples. We approximate the MTJ’s gradient using a straight-through estimator (STE), clamping values outside its saturation range to avoid exploding gradients. The MTJ can then be connected to the layer’s weights via:
Where weight quantization, array-splitting, and normalization algorithms are considered with respect to the layer’s weights. For each subarray , the layer output is the sum of all terms. As a result, for a given -long input stream and weight slices, terms would exist. In the final MVM output, these PS terms are scaled (through shifting) differently based on their and indices, such that a set of scalars can be defined. Each of these scalar terms is applied to an MTJ-sampled PS, making the partial derivative a set of scalars. Our MTJ in backpropagation is treated as a straight-through estimator (STE), and the MVM output of each PS array results from well-defined linear operations. The weight quantization’s gradients are similar to that of [16]. The gradient of batch normalization operation is also a scalar. Considering all of the operations involved in generating the MVM output, the gradient representation for backward propagation can be reduced to:
Fig. 4 compares the data distribution of crossbar output using the stochasticity-aware trained network (StoX) and the model trained with deterministic 1b PS (SA). We observe that incorporating stochasticity in training generates a broader and less polarized distribution with more intermediate values. Conceptually, such a distribution will facilitate using more of the sloped region of the tanh curve of switching probability. In contrast, the training with deterministic 1-bit SA will encourage the PS output to be concentrated around 0 and 1.
3.2.3 Error Mitigation with Multi-sampling
To mitigate the errors due to stochastic bit generation, we propose multi-sampling to recover accuracy with a trade-off of hardware efficiency (latency and energy). Counting from more SOT-MTJ samples will lead to a multi-level representation of the analog output. We explored up to 8 samples per conversion since the total conversion energy and latency increase linearly with increasing number of samples. To only increase the number of samplings for the layers that are most sensitive to PS precision, we perform a Monte Carlo simulation on our network’s trainable layers to guide an inhomogeneous multi-sampling scheme. For each layer, we apply a uniform random perturbation to its weights at inference and then measure the significance of the weights of that layer to the model accuracy. Fig. 5 shows the first layer being most susceptible to perturbation, indicating the most significance, while layers close to the output classifications are only moderately susceptible to error. Based on this analysis, we implement a model with mixed layer-wise sampling numbers (“Mix”), where layers with higher (lower) sensitivity are given more (fewer) MTJ samples. We achieve an accuracy similar to a homogeneous 4-sample network with only a small increase in operations compared to the 1-sample network (see discussions on the result in Section 4.3).
4 Experimental Results: Functional Accuracy and Hardware Efficiency
4.1 Evaluation Methodology
We evaluate both the functional accuracy and hardware efficiency of StoX-Net designs based on implementing convolutional neural networks for image classification on the MNIST and CIFAR10 datasets. Our evaluations include both a quantized first convolution layer (QF) and a high-precision first convolution layer (HPF). All subsequent layers in the evaluated models adopt the proposed MTJ-based stochastic conversion. All QF models take 8 samples per MTJ conversion in the first layer due to the importance of Layer 1 as shown by the Monte Carlo analysis in Fig. 5. When deterministic 1-bit sense amplifiers (SA) are used for PS processing, we label it as “HPF+1b-SA”.
We denote X-bit weight, Y-bit activation, and Z-bits per slice as . For example, a software-level model with 4-bit activation and 4-bit weight mapped to 2-bits per slice is . The baseline StoX network has , , , and 1 sample per MTJ conversion is applied using HPF configuration. The naming of the evaluated models highlights their differences from this baseline. The integer in front of the abbreviated letters (such as QF and HPF) means the number of MTJ sampling for all the other layers except the first layer conv-1. For example, 1-QF (4-QF) uses 1 (4) stochastic sample(s) of the MTJ converters in all layers except the quantized first layer, which still requires 8 samples per MTJ conversion. Mix-QF uses mixed sampling rates based on the Monte Carlo simulation with an 8-sample quantized first layer.
As the PS processing of the first layer and later layers can be set separately (see discussions on Fig.7),“1b-SA, 1b-SA QF” means that step-like 1b-SA (tanh with high ) is applied to all layers (including the first conv layer), showing the lowest accuracy due to severe quantization loss. “1b-SA, QF” applies stochastic MTJ with 8 samples to the first layer while 1b-SA is applied to all other layers. “1b-SA, HPF” is our version of the 1b deterministic “HPF+1b-SA” in the reference. We also include a model with binarized weight and activation (1w1a) to evaluate the performance of our approach on extremely quantized models. “Unsliced” refers to the case where the number of bits per memory device in hardware is the same as the number of bits in a weight element, while “Sliced” has 1 bit per memory cell.
@@ 4.1 Evaluation Methodology
We train ResNet-20 models based on the proposed convolutional methodology in Section 3. For CIFAR-10, we train with hyperparameters similar to [16], and for MNIST we lower epochs from 400 to 25. We note that larger models and more channels (such as such as VGG-9 or ResNet-18) can achieve higher accuracy for the evaluated benchmark tasks. However, our development focuses on ResNet-20, which has a compact model size and a small number of channels. We compare our StoX-Net results with similar implementations from the literature.
Moreover, we compare the hardware efficiency of the implementation of StoX-Net architecture with a standard IMC, with standard high-precision ADC for all layers (HPFA). Moreover, a sparse configuration (SFA) with reduced (full precision - 1)-bit ADC is also included as a stronger baseline for comparison. We evaluate hardware efficiency by simulating a standard IMC architecture similar to that of ISAAC [1], as shown in Fig. 6. We model and evaluate our architecture using Accelergy/Timeloop [22, 23, 24]. Our baseline HPFA architecture results were verified to be consistent with previous works [25][26]. Signed weights are represented using 2 cells per weight as in [6], leading to the possibility of having both positive and negative currents in the column bit lines. We use Successive Approximation Register (SAR) ADC with the area and power metrics scaled for 28nm technology based on comprehensive surveys and analysis of ADC architectures [7] [25, 27]. Table II summarizes key parameters for the hardware components. Because ADCs dominate the resources of IMC architectures in terms of energy, latency, and area [28], we focus on the impact of replacing the ADC with the stochastic MTJ-converter.
| Component | Energy/Action | Area/instance |
| DAC [22] | 2.99e-2 | 0.127 |
| Xbar Cell(1b/2b) [26] | (6.16e-3/4.16e-3) | 0.0308 |
| ADC(FP/Sparse) [7] | (2.137/1.171) | (6600/2700) |
| MTJ-Converter | 6.14e-3 | 1.47 |
| StoX-Net | HPF+1b-SA [6] | |||
| Samples | 1-QF | 4-QF | Mix-QF | / |
@@
4.2 Functional Accuracy
The proposed StoX-Net can reach within 1% of the reference HPF model’s accuracy for both MNIST and CIFAR-10 datasets, as summarized in Tables III & IV. Fig. 7 presents the ablation study of our algorithmic approach, The impacts of various techniques are summarized in panel (E). Both (A) and (E) panels of Fig. 7 show that applying 1b PS quantization to all layers (“1b-SA, 1b-SA QF”) suffers a significant loss, which is consistent with the state of the art that requires a high-precision first layer. Replacing 1b-SA with a multi-sample stochastic MTJ improves the accuracy, and using the largest sampling numbers achieves the highest accuracy. By enabling an 8-sample stochastic conv-1 layer only, “1b-SA, QF” experiences immediate (over 3.4%) accuracy improvement over “1b-SA, 1b-SA QF”. Among the HPF configurations, while “1b-SA, HPF” is slightly better than the 1-HPF StoX baseline, more samplings per MTJ clearly improve the accuracy of StoX model. Such observations suggest that multi-sampling in stochastic PS processing is an effective knob to enhance the accuracy of implementing IMC with aggressive PS quantization.
Fig. 7 (A) also shows that large arrays suffer accuracy degradation as more information loss occurs when PS quantization is applied to an increased range of possible crossbar output values. Fig. 7(B) demonstrates that stochastic multi-sampling consistently mitigates accuracy loss. In Fig. 7(C), bit slicing helps mitigate the accuracy degradation of quantization. Fig. 7 (D) shows that for models with one sample, a smaller (a more sloped tanh curve) is more erroneous due to stochasticity-induced fluctuation, while high (step-like) is more deterministic and less noisy. However, we re-iterate that with a sloped tanh curve (small ), better accuracy can be obtained through multi-sampling (Table IV). Such accuracy improvement is not possible with deterministic 1b-SA. Interestingly, the proposed “Mixed-HPF” manages to only increase the number of MVM conversions by 14.3% compared to the 1-sample model, but reaches 85.5% accuracy (similar to the 4-sample network’s performance). As discussed in the following section, a small increase of samples per MTJ conversion achieves the best trade-off in terms of achieving high hardware efficiency and maintaining accuracy.
4.3 Hardware Efficiency
@@
The proposed area-efficient and highly parallel PS processing directly impacts the datapath design pattern of crossbar MVM operation. Fig. 8 shows the pipelined operations inside a crossbar using an ADC (top) and a stochastic MTJ converter (bottom). In standard IMC, the length of each pipeline stage is determined by the longest stage, i.e. the ADC readout of all shared columns in a crossbar. Our optimization alleviates the throughput bottleneck by parallelizing all columns with compact MTJ converters, thereby considerably reducing the length of the pipeline stage.
@@
Our hardware performance evaluation is summarized in Fig. 9. Full precision ADC (HPFA) and Sparse design (SFA) are our baselines for comparison. The first observation is that SOT-MTJ converters significantly enhance parallelism with reduced area overhead, showing large savings in both area and EDP. Additionally, our design utilizes weights and activations of reduced precision, contributing further to efficiency. Processing reduced precision models with our proposed hardware design achieved improvements of area, energy, and latency by up to x, x, and x, respectively. Our design achieves up to x (x) EDP improvement compared to an IMC with sparse (full-precision) ADC.
We further evaluate the trade-off of efficiency and accuracy by varying the number of stochastic samplings. Multi-sampling improves accuracy compared to 1-QF, as observed in Fig.7, at the cost of reduced throughput and more energy consumption. Searching for the optimal trade-off needs to minimize the increase of samplings while achieving high inference accuracy. By exploiting the Monte-Carlo-based error sensitivity analysis shown in Fig. 5, we managed to only implement 2 or 4 samplings to a few layers in the deep neural network while maintaining the majority of PS processing at 1 sampling. As a result, such Mix-QF models only slightly increase the total number of MTJ conversions compared to the 1-sample models, maintaining relatively high hardware efficiency (x EDP gain) while achieving accuracy close to the 4-sample network (see Table IV). We further demonstrate in Fig. 9(b) that our architecture can be scaled to larger models and workloads while maintaining remarkable efficiency improvement.
5 Conclusion
@@
We develop StoX-Net, an efficient IMC design that eliminates the ADC bottleneck by processing array-level partial sum using stochastic MTJ. Based on our hardware-software co-design with stochasticity-aware training, DNN inference accuracy reaches close to the state-of-the-art accuracy while achieving up to 10x area efficiency and 24-130x improvement in hardware efficiency characterized by EDP. Moreover, we demonstrate that leveraging the multi-sampling capability of stochastic PS can quantize the first convolution layer (within accuracy loss compared to other quantized models), which is not possible in prior quantization-aware training. Future work will explore applying stochastic architectures to larger DNNs and datasets. More sophisticated designs of DNN architecture beyond standard networks (such as ResNet) can be created using automated machine learning such as network architecture search. Our work suggests exciting opportunities for utilizing stochastic computation and spintronics to develop next-generation AI hardware accelerators.